OpenStudy (anonymous):

May you check my answers? I've heard mixed views for 9 ; I got A. 10. D @mathstudent55

OpenStudy (anonymous):

First one is commutative property of multiplication

OpenStudy (anonymous):

i'm pretty sure the second one is associative property of addition

OpenStudy (mathstudent55):

9 is associative.

OpenStudy (mathstudent55):

10 is distributive

OpenStudy (anonymous):

Yay! :D

OpenStudy (anonymous):

May you check this one? I got commutative @mathstudent55

OpenStudy (anonymous):

commutative , I got A.

OpenStudy (mathstudent55):

I think step 1 is associative.

OpenStudy (mathstudent55):

Maybe step 1 is commutative. Maybe the correct answer is A.

OpenStudy (anonymous):

Woah... I was talking about 11...

OpenStudy (mathstudent55):

Yes, for 11, I think A is correct.

OpenStudy (anonymous):

10 looks like this: 4+6 = 2(2+3) 10 =2(5) 10= 10 @mathstudent55 So I got D..

OpenStudy (mathstudent55):

10 is definitely D. You are correct.